Victoria University Antarctic Research Expedition Science and Logistics Reports 1980-81: VUWAE 25
We were ready to start up Erebus on 4 December, but the weather was unsympathetic to us right from the start.
Between 9 and 19 December, Kienle, Estes and Kaminuma installed the telemetry receiver and recorder at Scott Base, and the geophones and transmitter at Abbotts Peak and Hoopers Shoulder on the flanks of Erebus. The main party of 10 including the writer, flew to the acclimatisation camp at the Fang on 10 December, and 8 of us reached the summit by helo on 17 December.
The second party flew to the Fang on 19 December. Continuous cloud developed there (although the summit remained clear) and they walked up a few at a time, the last arriving at the Summit Hut on 24 December.
In all, 105 man days were spent acclimatising instead of the planned 30.
By 19 December I had installed a low frequency microphone (8 inch Philips Hi Z speaker in 2 cubic foot sealed enclosure) and pre-amplifier in position on the rim between the main and side craters. It was connected by cable to an amplifier and monitor speaker in the Hut. The frequencies below 50 Hz were also recorded on channel 3 of tape seismograph A. Channels 1 and 2 were connected to a Willmore seismometer (Mk 1, Z, T0 = 1s) installed 220m towards the crater from the Hut.
By 20 December, Terai and Osada were recording a 4 geophone array (of 2km diameter surrounding the Hut) on their 2 tape seismographs.
By 23 December a figure of 8 induction loop, consisting of 3.6km of single conductor 1.5mm2 PVC appliance wire, had been laid around the main crater rim and the Hut area by Prosser, Parish, Summerville, Terai and myself.
The crater rim loop (3) was recording on channels 1 (high gain) and 2 (low gain) of tape seismograph B, and the Hut loop (4) on channel 3 (separate high gain amplifier) as shown in Figure 8.
By 26 December the loops were completely buried c. 50mm deep, eliminating noise due to the wire moving in the wind, but static discharge spikes limited the usable gain until bypassed by 1000 microfarad capacitors across the loops on New Year's Day.
Scott Base were asked to tune the spare discriminator to 1700 Hz and to monitor the rectifie microphone signal. Subsequently they reported an extremely high level of signal.
On 4 January signals from the microphone were reduced by connecting a 10 ohm resistor across the microphone cable at the VCO box. The modulation of the 1020 Hz sub-carrier (and presumably the 1700 Hz one also) was then peaking at 10% in a 15 knot wind. This was determined at the Summit Hut by means of the Alaska Test Receiver and a 1020 Hz discriminator, and of course an anemometer.
On 5 January the loops were disconnected from tape seismograph B and connected in figure of 8 configuration to the 1360 Hz VCO and transmitter. All signals were then being telemetered, as in Appendix 2C, but our communication radios had failed and Scott Base could not be asked to monitor the signals and report the background level and number and size of events back to us, so that we could optimise the adjustments, and calibrate the system.
The entire party returned to Scott Base on 9 January. The Abbott Peak, Hooper's Shoulder, and Summit geophones were recording correctly, but of my three channels only the 1700 Hz microphone channel was connected, and it was tuned to 3700 Hz instead of 1700 Hz. Generous help from Tom Earle and Stan Whitfield enabled us to get the 1020, 1360 and 1700 Hz discriminators installed and the channels recording as in Appendix 2C (although the 1700 Hz discriminator was noisy due to a faulty component we could not find), and also my cargon packed before returning to New Zealand on 10 January (the last flight in time for my daughter's wedding).
Regrettably, there was insufficient time in which to confirm that good data from the microphone and the loop were being recorded. As expected, the signal levels on a quiet volcano during a calm day were very low - perhaps too low! The weather which trapped the entire party at the summit, and the lack of communications with Scott Base denied us the opportunity to get it right with certainty. Plate VIA, B and C show equipment at the transmitter, the Summit Hut and Scott Base.
The party had a compak radio and two VHF hand radios from NZARP, a USARP radio, three Japanese hand radios, two University hand radios, and two Lands & Survey pack sets.
The Compak was not air-lifted from Fang camp as intended, and was finally brought up on foot on 8 January.
The USARP set distorted our transmission so badly that we were never understood.
The Japanese, V.U.W. and Lands & Survey sets gave good local communication only.
C. Erebus telemetry recording equipment in the Science Laboratory at Scott Base. Mr. Takanami (left) points to the data recorder. MT. Estes (right) points to the telemetry discriminators below.
When the batteries were nearly all discharged (and the weather still unflyable) we requested permission to recharge them from our several accumulators. This was denied, and we lost contact with Scott Base at the very time we were completing the telemetry transmitters, and needed to know what was being recorded at Scott Base so that we could optimise the VCO amplifier gains.
The field assistant who had signed the receipt for the VHF radios would not agree to the electronics experts in the party connecting other batteries to them via the exposed terminals en the outside of the case. Days later, he admitted that he believed he was personally liable for any damage to the radio which might result, and would not accept the risk.
Possibly the wording on the receipt form should be changed so that personal financial risk does not overshadow the risk to the whole project incurred through lack of communications.
As seen from Scott Base and New Harbour, the weather at Erebus summit appeared to be good from 22 November to 3 December 1980. The Fang was in cloud and mist most days between 12 and 24 December. When helicopters could get in they came bearing vapour trails.
The summit area was unusually free of snow when we .arrived there on 17 December, testifying the previous good weather. Fog and snow-fall began on 23 December and from 25 December on, fog and blowing snow became general as shown in Appendix VI. It appears that the horizontal wind shear boundary, often evidenced by the Erebus plume rising vertically for a short distance and then spreading out horizontally, had descended below the summit level and was creating an orographic cloud cap.
The deterioration in the weather experienced on Erebus after Christmas in 1974/5 and 1978/9 may have been a similar phenomenon.
In my opinion, the best four week period on Erebus is usually 24 November to 22 December.
Our event was very dependent on Helo support. We had 7 moves and especially during the latter part of our field time some set backs were caused to our programme through delays in scheduled helicopter movements.
Our first put-in, to Portal Mt. was close to maximum helicopter range with a full pay load. Nevertheless it was accomplished with little difficulty apart from a slight navigation problem due to our being off the helicopter grid maps. Although no subsequent difficulty with navigation was encountered, we prepared maps with the entire route on one sheet, which we could hand to the helicopter pilot. Our two moves at long range were carried out with two helicopters; Portal Mt. - Alligator peak and Alligator peak - Mt. Kempe. We had suspected they might do this and had budgeted Helo hours accordingly.page 36
From Table Mountain onwards, (5th December onward) weather had deteriorated, so that we had cloudier conditions. A fairly stable pattern did develop, so that it was generally clear in the morning, clouding over about midday. Accordingly, we requested our helo moves for the early morning, but Helo Ops did not appear to take these factors into account since all our moves and attempts at moves were made either in the evening shift or late in the day. Both Hugh Webb (DOIC) and Roger Clark (OIC) were aware of this situation and we are grateful for their efforts on our behalf. This is particularly the case for two incidents - The first was on our move Knobhead - Mt. Crean, when we were picked up together with K32 at 1900 on 15th December, but could not proceed beyond Depot Nunatak because of low cloud over Skelton Nevee. They really tried, but had to abandon the attempt and deposit us at L. Fryxell. A lift was arranged for us early the next morning so that we were at Mt. Crean in the morning and that day had completed over 1/3 of our sampling. This was fortunate as next day the weather deteriorated and we were only just able to get all samples before a blizzard confined us to tents for four days.
Our final move from Mt. Crean - Scott Base also showed some misunderstanding. The intention was to fly us out with one helo making a shuttle flight. It was almost midday when the first flight arrived. It had four crewmen and 7 seats, so under the rules operating could only lift out three passengers, leaving 5 for the following flight. In addition, the weather was rapidly deteriorating, with a cloud bank hovering 1km away on the plateau. Mainly by strenuous intercessions on our behalf by the OIC and DOIC, the helo was persuaded to drop us and a crewman Marble Pt., reorganise its seating and return immediately to collect the remainder of the party. The plan was successful and we all returned to Scott Base by about 1800.
B. MAIN AREAS COVERED
The specific objectives were to sample rocks from two sets of red beds, regarded as overbank and lake deposits in an alluvial plain. The youngest, the Aztec siltstone is Upper Devonian in age and we sampled at reported suitably exposed outcrops. These were all at the head of the Skelton Nevee, at Portal Mt., Alligator Peak and Mt. Crean. At Portal Mt. we camped on the Nevee at the foot of the Eastern side of the slope and about 100m below the rock exposures. The site was easy to land at and reasonably sheltered on drift snow.
The Alligator Peak site to the South was up a gully immediately at the foot of the rock exposure. It provided a confined landing site and a rather confined campsite on a patch of snow on the side of a slope. Our proximity to the rocks paid dividends as even with our limited time we were able to collect a full set of very friable rocks.
The second set of red lake bed deposits were in the Terra cotta siltstone of Lower Devonian age. The best outcrops are at Mt. Kempe, Table Mt. and Knobhead.
The Mt. Kempe site at an elevation of 2,600m was on the Kempe Glacier, overlooking the Koetlitz G1. It was a magnificent site with a view to Ross I. and beyond. We could regularly see Beaufort I, 230km distant. We were about 3/4 hr walk from the rock sections up the N.B. ridge of the mountain.page 37
The Table Mt. site was on a small snow patch in a valley looking out to the junction of the Ferrar and Taylor Glaciers. The rocks were a 20 rain walk up a hillside and were exposed mainly under two overhangs.
The Knobhead site was in a rocky valley between Knobhead and Mt. Handsley. These rock exposures were the best and most complete of the Terra Cotta. They showed evidence of extensive burrowing and were more purple than those at Table Mt. We sampled the complete section in two different places. Our camp was in the valley, about 200m below the section. There is a large flat platform on the E. ridge level with the top of the section. It has a commanding view of the Taylor and Ferrar Glaciers and of Mt. Lister in the Royal Society Range. It would make an ideal campsite.
The site at Mt. Crean was on the south shoulder of the mountain. It was almost level with the Polar Plateau to the W. but fell away sharply to the Skelton Nevee on the East. Geologically it was in the Weller Coal Measures of Permian age. We took over the site vacated by Event K14. The Aztec red beds were on the southern buttress of the ridge, about 3/4 hr walk, where we obtained samples throughout the 100m extent of the Aztec exposure. The sun got onto the face at about 1900 hrs but owing to restraints of time and weather most of the sampling was undertaken in the morning or cloudy conditions without the sun. We looked directly down to the Portal, one of the access routes from the Skelton Nevee to the Polar Plateau.
During my second visit to Antarctica this season with Event K9, whilst awaiting the departure of the Benjamin Bowring, I was able to obtain more samples from the following areas:
Wright Valley. Samples were obtained from about 12 lamprophyre dykes outcropping on the south side of Lake Vanda. They are intruded into the granite basement and are believed to be of Ordovician age.
Miers Valley. We landed on the north side of Miers Valley, on a tableland dotted with small tarns, which at that time of year were ice-free and circled by algae. We sampled the numerous dykes which criss-crossed the region and again are believed to be of Ordovician age. Our pick up helo came earlier than scheduled, but we still managed to collect good samples from 12 sites.
(See Appendix VI and Transport sections).
All our sites were at relatively high altitude and we had very little difficulty with communications. The only time was during the blizzard at Mt. Crean, Saturday 20th and Sunday 21st December.
E. LOSS/DAMAGE TO EQUIPMENT
The only trouble we had was with the antifreeze dispenser on our drilling equipment. Despite the care taken with draining the water from the system before entering the field, when we started operating at Portal Mt., the pipes and hand gun were frozen. We thawed the system out with the motor exhaust and primus and flushed through with pure antifreeze. The piston seals on one handpiece were damaged, but after replacement the equipment gave little further trouble.
The calf-length boots with green soles issued to Stephen were very slippery and the soles cracked, so that another pair had to be sent in. The stitching also failed on his used muklaks, requiring replacement.
Our glove combination proved more successful than on VUWAE 23. We used woollen finger gloves with an outer lightweight industrial mesh glove woven from a very tough synthetic thread. They stood up very well to handling the samples, and our hands although becoming a little damp from the antifreeze mixture kept reasonably warm by donning overmitts whenever possible.
We had the customary small amount of trouble with primus seals, but one has learned to be wary of this problem.
The VUWAE outer clothing and boots were well used. With the continuously cool conditions we experienced, the older generation clothing caused some inconvenience. The boot problem has been mentioned. Our overtrousers were very still and cumbersome and could not be put on over our boots. We had to anticipate weather conditions before venturing out and for our work, mobility, which was lost with the overtrousers, was a great asset. Our duvets were sufficiently warm but the toggles and zips were either broken or broke in use. Although adequately clothed we were not the most suitably dressed (in contrast to our DSIR field assistant). For high altitude field work, some effort should be made to equip the personnel with up-to-date and relatively new clothing.
Three 4-stroke Snotric motor toboggans (NZARP Nos. 039, 038 and 016 - a single geared older machine) were used on the sea ice in McMurdo Sound by Event 14. Sno Trac 35 primarily assigned to the Lands & Survey surveyors (K3) was also used for a time by this event. From 26 November to 3 December Snotric 016 and Sno Trac 35 were used off Butter Point by the seismic survey party (Dibble and lies), while the remainder of K14 continued the sediment sampling programme in New Harbour with snotric's 039 and 038. For each vehicle a summary of the mechanical problems, repairs and regular maintenance executed page 39 by K14 are presented below. A Vehicle Itinerary is presented in Appendix V.
During the sea ice sampling programme the helicopter hours used, arranged on an opportunity basis, were mainly for fuel and parts resupply. A helicopter was scheduled for K14's move to Mt. Crean (Lashly Mts.) on 11 December; the return to Scott Base executed on 16 December.
Snotric 016: Carburettor gasket replaced after spraying petrol from faulty gasket (21 October).
Nearly one litre of oil added to crankcase (30 October), burning quite a lot of oil (estimated 1 litre per 100km).
The carburettor air intake disk was lost and replaced with a camera lens cap during the trip to Butter Point (7 November). The head manifold [unclear: of] sheared off (11 November). A new elbow and carburettor disk were fitted after being flown from Scott Base (12 November), the air filter wired on and 0.5 litres of oil added to crankcase. This finished the one litre synarctic oil field pack with no likelihood of receiving more from Scott Base which had also run out.
The spring anchorage bolt on the front ski sheared and was replaced near Cape Chocolate (14 November).
The front ski quide rail was rewelded and turned 180° (cf. 039), a new spark plug fitted and 1 litre of oil put into the crankcase at Scott Base (17 and 18 November). Also the small inadequate manifold guard was removed and replaced by a larger version which was wired on. This was done to stop the burning of holes in the operator's clothing. The wired-on air filter cover was completely removed during this period of maintenance.
On 21 November the trip from Scott Base to Butter Point was aborted because 016 broke down 20km from McMurdo Station. It could be started by blowing into the tank to get fuel into the fuel pump but would die again almost immediately. The pump and carburettor were removed and the party returned to Scott with 037 and 038 which ran very well with only two passengers.
A new fuel pump and cleaned carburettor were fitted. Fuel was now found to be varporising in the carburettor after a kilometer or so when the engine became warm. The cowling vents were removed to give better air circulation. 016 ran continuously thereafter but readjusting the carburettor did not regain the power the machine had earlier in the season when temperatures were cooler.
During the seismic programme one of the sprocket wheels on the rear axle sheared from the axle destroying a rear bearing assembly. A spare rear axle and bearing assembly were fitted. During this period the petrol tank was accidentally run dry. After refilling the machine would not start until the sump was filled with oil to give sufficient impulse for the petrol pump to operate. The front ski guide rail also sheared again leaving 016 without good steerage on the return to Scott (4 December).
Snotric 039: On 22 October this toboggan refused to start when returning to page 40 Scott Base from the desalinisation plant at McMurdo. A small spark was being generated but the engine did not fire when primed with raw petrol or ether. The machine was towed back to Scott where the condenser was replaced next day. A new fuel pump was fitted at the same time. The two air filter brackets broke as on 038 when returning from Evans to Scott (28-29 October).
All three machines were topped up with oil at Scott Base (30 October) and air filter brackets rewelded.
Oil topped up, 0.25 litres (12 November).
Air filter brackets broke again and manifold broke letting exhaust gas escape (warm knee), similar to 016 (15 November). Major repairs at Scott (17-18 November). New manifold made up (no spares at Scott). New single air filter brackets were made by bending strap steel (1" × 3/16") 270° around the filter pipe and welding, then bolting the bracket to the head with two head bolts. Brackets were made for both 039 and 038. The broken front ski guide rail was rewelded, turned 180° and refitted to the ski. New spark plug fitted and 0.5 litres of oil added to the gearbox. A new variator drive belt fitted (26 November) manifold broken once again (27 November). A new manifold fitting made up at Scott Base was fitted the next day (28 November). Oil topped up also and plug checked.
This machine then ran well up until 6 December when it was relinquished to K2.
Snotric 038: During the trip from Scott Base to Evans (24 October) the two brackets (1/2" × 1/8" steel) attaching the air-filter to the engine block snapped. The filter was then wired to the engine block but failed within five minutes of travelling. The fuel tank was accidentally run dry from Evans to Royds. Great difficulty was experienced to restart the engine (0.5 hour) and afterwards all machines were topped up regularly to prevent running dry.
Oil topped up and air filter brackets repaired (30 October) at Scott Base. In the field this machine had already become difficult to start, however, it was found to start better in the warmth of the Scott Base garage.
Oil topped up, 0.25 litres (12 November).
Snotric 038 had become increasingly hard to start (10-20 minutes pulling was not uncommon) cold and was also difficult to start hot, requiring ether on several occasions when hot. The machine was adjusted to run warm and once started did run reasonably well.
Air filter brackets broke once again when returning to Scott (15 November) and the spark plug was replaced making the cold starting slightly easier.
At Scott Base (17-18 November) the new airfilter strap was fitted (cf. 039), fuel pump replaced and the fuel line from the tank refitted to stop a primer bulb leak. Topped up crankcase (0.25 litres) and gearbox (0.5 litres) with oil. The new plug and fuel pump made cold starting considerably easier, although the temperatures were much milder than in October. Regular maintenance; oil, plug checked and a new variator drive belt fitted on 28 November. This machine also continued to run well for the remaining time in the field.page 41
Sno Trac 35: Prior to 25 November four new bogies had been fitted to 35. Two of these new bogies from the front right bogey assembly subsequently broke and were replaced enroute to Butter Point (25 November).
A steering clutch splined shaft was stripped during the seismic survey programme. The Scott Base Mechanic (Bruce Scott) flew to Butter Point where he replaced the steering clutch assembly on the sea ice.
Two further bogies were broken on the return to Scott Base from Butter Point (4 December).
General Summary: The three toboggans were in good condition when they were handed over to K14 on 14 October. This was due to the work done in early October by Nevil Clark and Bruce Scott (Base Mechanic). The assistance both these people gave to this event especially when initially testing the toboggans and then later in the season, was invaluable.
Nearly 1,000 kilometres were travelled by all three toboggans on relatively smooth often bare hard sea ice. The hard conditions were probably responsible for only some of the damage incurred, particularly the sheared axle (016) and steering ski guide rail breakages (016, 039). The other damage and mechanical failures which occurred were related primarily to the Briggs and Stratton engine. The vibration from these engines were responsible for the failure of the air filter fittings (once 016, twice 039, twice 038). This recurrent problem was only solved by building an extremely heavy fitting fastened by two head bolts but attached from only one place on the head. This prevented any opposing vibration which may have caused the two bracket mounting to fail. Manifold failure occurred on three separate occasions and can also be attributed to the engine vibration. The right angle elbow on 016 vibrated sufficiently to distort the male thread entering the block, eventually causing it to shear off. The manifold on 039; a threaded nipple pipe welded to a flange broke twice, in one case leaving part of the sheared nipple pipe in the exhaust port.
The carburettor was adjusted to give the best performance when running warm (after about 1/4 hour from a cold start early in the season). The idling speed was adjusted to prevent the machine stalling even when hot but meant the idling speed was too high for the Salisbury clutch to disengage when changing gears. However, by turning off the engine, changing gears then quickly turning on the engine it was possible to change gears relatively smoothly.
It was always necessary to have at least 2 metres of "slack rope" before being able to start moving a heavy sledge (=900 lbs) on the flat ice. Snow cover or a slight upward gradient made things even more difficult. The Briggs and Stratton engines have a particularly slow pick up so they were developing less than half revs by the time load came on the sledge. At these revs less than 10 BHP was being produced. If the sledge didn't move the moderate torque caused "track spin" and the toboggan "dug in" and made the situation more difficult. The rope starter makes starting these big engines even more difficult. Even when the engine was warm some of the event members could not start the machines consistently.page 42
Sno Trac 35 seems to specialise in destroying bogey wheels. These were all broken when travelling between 10 and 15km per hour on dead flat smooth bare sea ice, for distances of about 30km. The front bogey assemblies accounted for all the broken bogies which were found to be very hot at the time of breaking. This indicates that they were doing a lot more work than the absolutely cold bogies further aft.
Two dog sledges (B5 and B7) and a Tamworth (T1) were used initially to Cape Evans. The dog Sledges which did not have keels were replaced by a new Tamworth (T3) and a double-ended Nansen with keels (B4 on 1 November. The keeled sledge towed significantly straighter on hard snow than the previous dog sledges but was just as difficult to control on bare ice as the non-keeled sledges.
Loading: Both Tamworth sledges carried between 800 and 900 lbs for most of the time. This was probably too much for these sledges on bare sea ice, which is far more rigorous than snow. The alternative to distribute the load on tandam sledges was however impractical because; of the much greater difficulty to negotiate rough ice with two sledges, the limited manpower of our event, and the towing capabilities of the toboggans was often not sufficient to tow two moderately heavy sledges.
B4 was loaded to about 600 lb which was a suitable weight for sea ice conditions. The keels however may have been more effective if located further aft especially when close towing on bare ice.
Damage: T1 is an old sledge with non-laminated steamed timber bridges, and already had two cracked bridges on 15 October. The cracks became enlarged and the remaining bridges cracked during the programme. Hose clips provided a very satisfactory repair for the cracked bridges in the field.
T3 is a new laminated sledge which proved very satisfactory. No bridges were cracked but the laminated runners suffered some damage. Delamination occurred on both runners at all ends where the bridges are housed into the runners. This damage seemed to be a result of the design and could be prevented if the top lamination, which houses the bridges, were laminated further into the curve of the runners.
The proud lashing (thong leather) on the Tamworth sledges became worn from the movement of poorly-fitting sledge tanks. The lashings were replaced without much difficulty in the field.
Broken bridges, especially on an old sledge, worn lashings and towing bridles are all damages which should be expected as normal "wear and tear" in a season. The sea ice even though excellent travelling this season is extremely hard on this type of equipment.
B. MAIN AREAS COVERED
The bottom sampline programme undertaken by Event 14 required several traverses across the sea ice in McMurdo Sound. The areas of interest were from page 43 Scott Base to Cape Royds, across to Butter Point and in New Harbour and south to Cape Chocolate. A more detailed account is presented in the Itinerary (Appendix IV).
Difficult weather conditions restricted the Event 14 programme early in the season (25-29 October) when several southerly storms caused intermittent blizzard conditions in the McMurdo Sound area. Four days were spent tentbound during this period. "Normal" weather conditions for McMurdo Sound prevailed later in the season during the major period of the field programme. A summary of weather observations is presented in Appendix VI.
DSIR Compak SSB radios were used for field communications and worked satisfactorily. During the seismic programme two Compaks were used on a free channel (4700 KHz) between the recording station and shot firing point. The radios were also used to record the shot instant on the recording (see Appendix IF).
E. LOSS/DAMAGE TO EQUIPMENT
The major damage occurred to toboggans and sledges; detailed in the Transport section.
No equipment was lost or damaged. However, the sphincter corer, which was an important part of the operation, did not perform satisfactorily.
The sphincter corer was designed to take a shallow (=300mm long) large diameter (=200mm) minimally disturbed core and retain the sea floor sediment-water interface intact. The main feature of the corer is a cloth sleeve (sphincter) which cuts the sediment in situ and retains the core with a nearly water-tight seal.
Two major types of problems were encountered when operating the corer. A lack of penetration was evident in some areas of the Sound where the sea floor was stoney, e.g. near Ross Island. Thick sponge mat in some areas also contributed to poor corer penetration. Extra lead weights and a short barrel option will be used to overcome the penetration problem.
The second problem was that caused by freezing sea water in the trigger mechanism. This mechanism consisted of a main pin into which was hooked a sprung detent pin. Both pins were greased and enclosed in a close-fitting steel housing. Sea water on freezing ejects brine and forms "fresh water" ice. This ice does not melt in the sea water, which has a temperature of −1.9°C. Because the trigger was enclosed, ice was not flushed out as the corer descended to the bottom. The problem should be solved by opening out the trigger so that it will be self-draining, and can be flushed either in the sea or with alcohol.
In general the other scientific equipment performed well, particularly the winch. The salinity-temperature bridge was difficult to use in the open page 44 where it was impossible to keep unfrozen the water standards used for calibration of the instrument.
All transport for Event 15 was by helicopter and movements between camps was well-timed and efficient with only two delays due to bad weather. Considerable saving of helo time was afforded by good management by Scott Base D.O.I.C. who arranged moves so that flight times from Scott Base to the Dry Valleys was shared by several events moving on the same day.
B. MAIN AREAS COVERED
A single area between the Miers Valley and Salmon Valley from the coast to the Blue Glacier was covered. Five camps were established for the field party which enabled the 200km2 area to be mapped and samples in single day trips without having to establish secondary overnight camps.
|ii)||Upper Garwood valley at Collen Lake|
|iii)||Lower Garwood valley about 5km from the coast|
|iv)||Marshall valley about 5km from the coast|
Good weather conditions prevailed during Event 15 stay in the area between Miers valley and Salmon Valley. Weather observations were made and relayed daily to Scott Base. The unfavourable weather that was experienced did not hinder field work; snow falls usually covered outcrops for only a day or two. The weather became more unsettled and cooler in the latter part of the season (mid-December/early January).
Communication with Scott Base was maintained with a Labgear radio which operated effectively for the entire field trip. On only one day was it necessary to relay to Scott Base via Vanda Station due to ionispheric disturbance. Battery life was excellent due to low power being used for transmission on most days.
E. LOSS/DAMAGE TO EQUIPMENT
A mattress and a snow gaiter were lost in blustery conditions that were prevalent, particularly at the Salmon Valley campsite.